OpenStudy (anonymous):

1. A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air? (2 points) 2.Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart? (2 points)

OpenStudy (anonymous):

1. (velocity +40) *5 = distance (velocity -40) *6 = distance (velocity +40) *5 =(velocity -40) *6 see if you can do the algebra to solve for V (distribute, group like terms)